OpenStudy (anonymous):

what is the y-intercept of y=2/3x+2?

OpenStudy (anonymous):

To find the y intercept, you set x equal to 0

OpenStudy (anonymous):

Do you know what happens after you do that?

OpenStudy (anx):

Set x equal to zero y = 2/3(x) +2 Substitute 0 for x y = 2/3(0) +2 Multiply y = 0 + 2 Solve y = 2

OpenStudy (anonymous):